OpenStudy (anonymous):

a+1=√b+1 for b solve

OpenStudy (anonymous):

\[\sqrt{b}=a+1-1\] \[\sqrt{b}=a\] \[b=a^{2}\]

OpenStudy (anonymous):

a. b = a2+ 2a b. b = a + 2a2 c. b = 2a2+ a which one?

OpenStudy (callisto):

\[a+1=\sqrt{b+1}\]\[(a+1)^2=(\sqrt{b+1})^2\]\[(a+1)^2 = b+1\]\[b= (a+1)^2 -1 = ...?\] To expand (a+1)^2, you can use the identity (a+b)^2 = a^2 + 2ab+b^2

OpenStudy (anonymous):

ans.a a^2+2a

OpenStudy (anonymous):

\[b=a^{2}+2a+1-1\] \[b=a^{2}+2a\]

OpenStudy (anonymous):

@Luis_Rivera that is sqrt{b+1}.